# Let’s Take Esoteric Programming Languages Seriously

### v1

@@ 5.2.1. Comedy
Figure 5. Hypothetical esolang suggestion from the xkcd webcomic (number 2309)
Tiwari et al. (Tiwari et al., 2024) explore how humor is great motivation for software developers. Further, Kuutila et al. (Kuutila et al., 2024) analyse reddit posts to characterise programming humor. They show that ‘things that violate accepted patterns in a non-threatening way are perceived to be humorous’—attributes that clearly apply to esolang programs.
@@ 5.2.2. Challenge
For esoteric languages, programs need to be written differently, in an unconventional syntax and structure. Constaints are ever-present, and in some cases programs are practically impossible to write, e.g. for the Malbodge language. We discussed these issues earlier, in Section 3 when we were considering the attraction of writing code in an esolang. Similar challenges will apply to develop an esolang, whether or not that language is self-hosting.
@@ 5.2.3. Creativity
Finally, there is the attraction of creating a programming language ‘of one’s own’ (to misquote a phrase from Virginia Woolf). Constructing an esolang permits the programmer to enjoy the luxury of a ‘break from the bland, bloated multiparadigm languages with imperative roots’ (Kneusel, 2022).
This facility of expression, of ‘speaking code’ (Cox, 2013), is a powerful motivation for people to generate their own, custom esolangs. As we note in Section 4, creating an esolang gives excellent opportunity for students to reflect on what make a good programming language, and to identify tradeoffs in language design.
@@ 6. Reasons for Developing a New Programming Language
As we study the field of esoteric programming languages, we are confronted with a question that relates to more general programming language philosophy: Why does a person choose to design and develop a new programming language? Can we identify a set of overarching motivations for the construction of new programming languages?
In this section, we suggest five overarching themes pertaining to programming language design motivation:
@@ 6.1. Expressivity
Jones and Bonsignour note that a programmer can write a fairly constant number of lines of code per hour, regardless of the language used (Jones and Bonsignour, 2011). So a more expressive language should enable programmers to be more productive. They conclude that ‘the overall effort associated with coding and testing are much less significant for high-level …languages than for …low-level languages”.
@@ 6.3. Education
However, it is noteworthy that no mainstream industrial strength languages were explicitly designed for educational use, or with respect to pedagogical principles (Kölling, 2024). Further, initially simple industrial languages accumulate new features throughout their lifetime, generally measured in decades (Favre, 2005).
@@ 6.4. Economy
Some languages are commercial imperatives—needed to satisfy economic ends. The language that springs to mind most readily is C#, originally conceived ‘direct rival’ to the Java ecosystem, which seemed to be a ‘deep challenge to Microsoft’ (Syme, 2020).
@@ 6.6. Esolang Motivations
In terms of economy, esolangs are not commercially viable in any sense. No meaningful software is developed in esolangs and no esolang toolchain is available on a commercial basis. Perhaps the only people who might benefit financially from esolangs are book authors and publishers. For instance, MIT Press has a current esolang title available (Cox, 2013) and another one forthcoming (Temkin, 2025).
@@ 8. Related Work
Mateas and Montford (Mateas and Montfort, 2005) discuss the sociological and phenomenological issues around esolangs. They link this community with the International Obfuscated C Code Contest (IOCCC), which is a celebration of unconventional code expressed in a mainstream programming language (i.e. C) but where the syntax and semantics of the language are highly abused in order to generate code that looks unusual (some kind of ASCII art, generally) and does something unexpected when compiled (some kind of textual output, normally). Cox (Cox, 2013) explores various non-technical aspects of esolangs in further depth. Temkin (Temkin, 2017) gives the most complete review of the esolang culture and philosophy. He seeks to explore the nature of esolangs, and the motivation for their design and use. His findings broadly concur with ours. His more recent study (Temkin, 2023) posits that esolangs enable ‘personal expression and …elegance within chaos’ and that they ‘challenge base assumptions of who languages are designed for and how they should be used’. Esolangs favour individual personality over practicality, and expression over comprehension.
@@ 9. Conclusion
Ultimately, the use and development of esolangs is a form of freedom of expression (Cox, 2013). Such freedom must be preserved and valued by the community. Even if that freedom of expression allows people to express themselves in highly limited ways, the freedom to be so constrained must still be preserved.
@@
19. A language that doesn’t affect the way you think about programming, is not worth knowing.

### v2

@@ 5.2.1. Comedy
Figure 5. Hypothetical esolang suggestion from xkcd.com webcomic, number 2309 ($C$ BY-NC 2.5 Randall Munroe)
Tiwari et al. (Tiwari et al., 2024) explore how humor is great motivation for software developers. Further, Kuutila et al. (Kuutila et al., 2024) analyse reddit posts to characterize programming humor. They show that ‘things that violate accepted patterns in a non-threatening way are perceived to be humorous’—attributes that clearly apply to esolang programs.
@@ 5.2.2. Challenge
For esoteric languages, programs need to be written differently, in an unconventional syntax and structure. Constraints are ever-present, and in some cases programs are practically impossible to write, e.g. for the Malbodge language. We discussed these issues earlier, in Section 3 when we were considering the attraction of writing code in an esolang. Similar challenges will apply to develop an esolang, whether or not that language is self-hosting.
@@ 5.2.3. Creativity
Finally, there is the attraction of creating a programming language ‘of one’s own’. Constructing an esolang permits the programmer to enjoy the luxury of a ‘break from the bland, bloated multiparadigm languages with imperative roots’ (Kneusel, 2022).
This facility of expression, of ‘speaking code’ (Cox and McLean, 2012), is a powerful motivation for people to generate their own, custom esolangs. As we note in Section 4, creating an esolang gives excellent opportunity for students to reflect on what make a good programming language, and to identify tradeoffs in language design.
@@ 6. Reasons for Developing a New Programming Language
As we study the field of esoteric programming languages, we are confronted with a question that relates to more general programming language philosophy: Why does a person choose to design and develop a new programming language? Can we identify a set of overarching motivations for the construction of new programming languages? Here we are using our analysis of esolangs to reflect on wider aspects of programming language design.
In this section, we suggest five generic themes pertaining to programming language design motivation:
@@ 6.1. Expressivity
Jones and Bonsignour note that a programmer can write a fairly constant number of lines of code per hour, regardless of the language used (Jones and Bonsignour, 2011). So a more expressive language should enable programmers to be more productive. They conclude that ‘the overall effort associated with coding and testing are much less significant for high-level …languages than for …low-level languages’.
@@ 6.3. Education
However, it is noteworthy that no mainstream industrial strength languages [^9] One reviewer kindly pointed out that Pascal might be an exception here. were explicitly designed for educational use, or with respect to pedagogical principles (Kölling, 2024). Further, initially simple industrial languages accumulate new features throughout their lifetime, generally measured in decades (Favre, 2005).
@@ 6.4. Economy
Some languages are commercial imperatives—needed to satisfy economic ends. The language that springs to mind most readily is C#, originally conceived as a ‘direct rival’ to the Java ecosystem, which seemed to be a ‘deep challenge to Microsoft’ (Syme, 2020).
@@ 6.6. Esolang Motivations
In terms of economy, esolangs are not commercially viable in any sense. No meaningful software is developed in esolangs and no esolang toolchain is available on a commercial basis. Perhaps the only people who might benefit financially from esolangs are book authors and publishers. For instance, MIT Press has a current esolang title available (Cox and McLean, 2012) and another one forthcoming (Temkin, 2025).
@@ 8. Related Work
Mateas and Montford (Mateas and Montfort, 2005) discuss the sociological and phenomenological issues around esolangs. They link this community with the International Obfuscated C Code Contest (IOCCC), which is a celebration of unconventional code expressed in a mainstream programming language (i.e. C) but where the syntax and semantics of the language are highly abused in order to generate code that looks unusual (some kind of ASCII art, generally) and does something unexpected when compiled (some kind of textual output, normally). Cox (Cox and McLean, 2012) explores various non-technical aspects of esolangs in further depth. Temkin (Temkin, 2017) gives the most complete review of the esolang culture and philosophy. He seeks to explore the nature of esolangs, and the motivation for their design and use. His findings broadly concur with ours. His more recent study (Temkin, 2023) posits that esolangs enable ‘personal expression and …elegance within chaos’ and that they ‘challenge base assumptions of who languages are designed for and how they should be used’. Esolangs favour individual personality over practicality, and expression over comprehension.
@@ 9. Conclusion
Ultimately, the use and development of esolangs is a form of freedom of expression (Cox and McLean, 2012). Such freedom must be preserved and valued by the community. Even if that freedom of expression allows people to express themselves in highly limited ways, the freedom to be so constrained must still be preserved.
@@
Epigram 19: A language that doesn’t affect the way you think about programming, is not worth knowing.
@@
Acknowledgements.
We would like to thank the following people who gave us useful additional comments which improved this paper: Alan Blackwell, Norman Gray, and John Levine. This work was supported in part by UKRI/EPSRC under grant Sponsor EPSRC https://epsrc.ukri.org Grant #M4Secure.